A stone is falling freely from rest and the total distance covered by it in the last second of its motion equals the distance covered by it in the first 3 seconds of its motion. How long does the stone remain in air?


1
Expert's answer
2021-06-22T16:59:41-0400

Answer:-

"s =ut + 0.5gt^2" ........1

In first 3sec, let it covers distance "d"m.

"d=0.5g(3)^2 =4.5g"

This distance is same as distance traveled in last second

So, according to formula,

"Sn=u+0.5a(2n\u22121)"

"4.5g=0.5g(2n\u22121)"

Here, u=0 and n is number of seconds for which stone is in air.∴n=5 sec

Substituting in eqn 1,

"\boxed{S= 122.5m}"